OpenStudy (haleyelizabeth2017):

A merry-go-round ride has outer horse figures that go up and down exactly three times during one rotation of the carousel. One of the high-points for each horse occurs where a rider is just close enough to reach out and try to grab a metal ring from a mechanical dispenser. If the rider succeeds in grabbing a brass ring instead of an iron ring, the rider has won a free ride on the merry-go-round. Hence the expression, "reaching for the brass ring." Write a sinusoidal function that models the height of one of the horse figures as a function of the rotation of the main carousel with theta = 0 at the ring dispenser. The amplitude of the horses vertical motion is 1.1 meters around an average height of 1.6 meters.

OpenStudy (phi):

it sounds like we want to use cosine because cos(0) (at the "ring dispenser") is at a max and we want "One of the high-points for each horse occurs where a rider is just close enough to reach out and try to grab a metal ring from a mechanical dispenser"

OpenStudy (haleyelizabeth2017):

but we have to use sin...

OpenStudy (phi):

if the carousel rotates at theta, the horse is going up and down 3 times faster so 3theta so you want \[ A \cos (3 \theta) + B\] you need A and B if you have to use sin, shift the sin pi/2 to the left i.e use \[ A \sin\left(3 \theta + \frac{\pi}{2}\right) +B\]

OpenStudy (haleyelizabeth2017):

Ah okay

OpenStudy (haleyelizabeth2017):

\[1.1 \sin\left(3 \theta + \frac{\pi}{2}\right) +1.6\]How did you get \[3\theta ?\]

OpenStudy (haleyelizabeth2017):

Oh wait....I misread at first x'D

OpenStudy (phi):

outer horse figures that go up and down exactly three times during one rotation

OpenStudy (phi):

up and down means (to me) one full period

OpenStudy (haleyelizabeth2017):

Okay :)

OpenStudy (haleyelizabeth2017):

So if we had to use sin like this time...and it's similar to this, would it be \(\frac{\pi}{2}\) most of the time?

OpenStudy (phi):

in this problem, you want the sin to peak at theta =0 However, you (should) know that sin peaks at pi/2, so shift the sin to the left by pi/2 shift to the left means add pi/2 (shift to the right would be - pi/2) If they problem is to align the peak at say theta=pi/4 (45 degrees) rather than 0, you would shift the sin from pi/2 to pi/4 which means pi/4 to the left, i.e. add pi/4

OpenStudy (haleyelizabeth2017):

ah

OpenStudy (haleyelizabeth2017):

Thank you very much @phi :)